# Beyond Bradley-Terry Models: A General Preference Model for Language Model Alignment

### v2

@@ 2 Related Work
Figure 2: Intransitiveness in real-world preferences.
Reward-Based Reinforcement Learning from Human Feedback (RLHF).
Typical approaches to modeling human preference for language model alignment usually learn a reward model from a preference dataset. The human preference is assumed to follow the Bradley-Terry (BT) model (Bradley & Terry, 1952) or the Thurstone model (Thurstone, 2017). LLM policies then are fine-tuned to maximize these scalar reward signals for better alignment (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022). Later, the direct preference optimization (DPO) methods are proposed by Rafailov et al. (2024) only implicitly to learn a reward model represented by an LLM. The human preference is still assumed to follow the Bradley-Terry model. However, the reliance on scalar rewards imposes a total ordering on preferences, which may not reflect the intransitive or stochastic nature of human judgments (Tversky, 1969; Agranov & Ortoleva, 2017).
Preference-Based Reinforcement Learning from Human Feedback.
Recently, there emerged a line of works that directly estimates the preference probability without imposing a reward-based preference model or any transitivity assumptions (Lou et al., 2022; Wu et al., 2023; Wang et al., 2023) either for preference-based RL or in the context of RLHF. Efforts have been made to optimize policies directly from pair-wise preference comparisons, thereby mitigating the limitations of scalar reward functions (Munos et al., 2023; Swamy et al., 2024; Rosset et al., 2024; Wu et al., 2024b).
@@ 3.2 Preference Modeling
3.2.1 Supervised Pair Preference Models
@@ 3.3 Preference-based Reinforcement Learning from Human Feedback
$\max_{\pi}\ \mathbb{E}_{\mathbf{x}\sim\mathcal{X}}\bigg{[}$ $\mathbb{E}_{\mathbf{y}\sim\pi(\cdot\mid\mathbf{x}),\ \mathbf{y}^{\prime}\sim\mu(\cdot\mid\mathbf{x})}\Big{[}\mathbb{P}\big{(}\mathbf{y}\succ\mathbf{y}^{\prime}\mid\mathbf{x}\big{)}\Big{]}-\beta\,\mathrm{KL}\big{(}\pi\|\pi_{\text{ref}}\big{)}\bigg{]},$
@@ 4.2 Implementing General Preference Embedding Model
To effectively capture general preferences while maintaining computational efficiency, we implement our preference embedding model by augmenting an existing language model with two additional components: an eigenvalue scale gate and an eigenvector embedding head.
@@
where $\mathbf{D}(\mathbf{x})$ is a block-diagonal matrix with blocks $\sqrt{\lambda_{l}(\mathbf{x})}\mathbf{I}_{2}$, and $\mathbf{R}^{\succ}$ is the skew-symmetric preference operator. We normalize the embeddings $\mathbf{v}_{\mathbf{y}}$ to have unit length to ensure training stability.
@@ 5 General Preference Optimization
$\bm{\theta}_{t+1}=$ $\arg\min_{\bm{\theta}}\mathbb{E}_{\mathbf{x}\sim\mathcal{X},\mathbf{y}\sim\pi_{\bm{\theta}_{t}}(\cdot\mid\mathbf{x})}\Bigg{[}\bigg{(}\log\frac{\pi_{\bm{\theta}}(\mathbf{y}\mid\mathbf{x})}{\pi_{\bm{\theta}_{t}}(\mathbf{y}\mid\mathbf{x})}-\frac{1}{\beta}\big{(}\widehat{s}\left(\mathbf{y}\succ\pi_{\bm{\theta}_{t}}\mid\mathbf{x}\right)-\log Z_{\pi_{\bm{\theta}_{t}}}(\mathbf{x})\big{)}\bigg{)}^{2}\Bigg{]},$ (5.3)
@@
Consider the optimization problem defined by the GPO loss (5.3) and assume it is realizable. Let $\{\pi_{\bm{\theta}_{t}}\}_{t=1}^{T}$ denote the sequence of policies generated by GPO, and define $\bar{\pi}_{T}=\frac{1}{T}\sum_{t=1}^{T}\pi_{\bm{\theta}_{t}}$ as the average policy. Given that the preference score $s$ is bounded within $[-\rho,\rho]$, by setting $\beta=\Theta\left(\sqrt{T}\right)$, we have:
@@
$\mathbb{E}_{\mathbf{x}\sim\mathcal{X},\mathbf{y}\sim\pi_{\bm{\theta}}}\bigg{[}\bigg{(}\widehat{s}(\mathbf{y}\succ\pi_{\bm{\theta}_{t}})-\beta\log\frac{\pi_{\bm{\theta}}(\mathbf{y}|\mathbf{x})}{\pi_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})}\bigg{)}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(\mathbf{y}|\mathbf{x})\bigg{]}$
@@
So Equation (5.3) can also be seen as a policy gradient method for the optimization problem (5.1).
@@ 6 Experiments
Table 3: AlpacaEval 2.0 evaluation results. Base model: Llama3-8B-it, Evaluator: GPT-4-turbo. The results are grouped by the size and type of the RM or PM, and the number of iterations. Bold entries indicate that GPM outperforms BT RM under the same training settings.
@@ 6.1 Experiments on RewardBench
Results and Analysis. The results are presented in Table 2. On RewardBench, using the Gemma-2B-it base model, GPM achieves an average score of 82.29%, which is an improvement of 7.44% over the BT model’s average score of 74.85%. Specifically, in the Chat task, GPM improves performance from 67.32% (BT RM) to 79.61%, and in the Chat-Hard task, from 63.37% to 75.66%. For the Llama-3.1-8B-Instruct base model, GPM achieves an average score of 91.90% (embedding dimension 8), representing a 1.34% improvement over the BT model’s average score of 90.56%. In the Chat task, GPM improves from 88.55% (BT RM) to 93.58%, and in the Chat-Hard task, from 85.75% to 88.38%. Using the Gemma-2-9B-it base model, GPM achieves an average score of 92.33% (embedding dimension 4), showing an improvement of 0.87% over the BT model’s average score of 91.46%. In the Chat task, GPM boosts performance from 91.62% (BT RM) to 93.58%, and in Chat-Hard, from 85.96% to 87.72%. These results indicate that GPM outperforms the BT model across various base models and tasks, particularly in the Chat and Chat-Hard tasks. Note that BT RM is a special case of GPM when the embedding dimension $d=1$ (see Section 4.3).
@@ 6.3 Downstream Performance on Aligning Language Models with Human Preferences
We further investigate the effectiveness of GPM in language model for alignment using Self-Play Policy Optimization (SPPO) (Wu et al., 2024b) and our proposed General Preference Optimization (GPO), integrating preference scores provided by our GP representation model (GPM). We evaluated the models on AlpacaEval 2.0 (Dubois et al., 2024), MT-Bench (Zheng et al., 2023), GSM8K, MMLU, etc., several widely used benchmarks for evaluating LLM alignment.
Results and Analysis. The evaluation results on the benchmarks are as follows. For AlpacaEval 2.0, we compared the generated responses of the aligned models with those of GPT-4o-mini and GPT-4-turbo. The results of the three evaluators are presented in Tables 3 and 5. From Table 3, we observe that both SPPO and GPO demonstrate improved win rates with successive iterations, highlighting the iterative nature of these optimization methods, and GPO consistently outperforms SPPO. In addition, the bolded entries indicate that GPM-integrated methods consistently outperforms BT RM-based methods under the same settings on Win Rate.
@@ 2. Magnitude preserving:
First, since the preference score $s$ is bounded in $[-\rho,\rho]$, we can normalize it to $[0,1]$ by the transformation:
@@
By Theorem 1 in Freund & Schapire (1999), for any sequence of mixed policies $\mu_{1},\mu_{2},\ldots,\mu_{T}$, the sequence of policies $\pi_{1},\pi_{2},\ldots,\pi_{T}$ produced by GPO satisfies:
$\sum_{t=1}^{T}\widetilde{s}(\pi_{t}\prec\mu_{t})\leq\min_{\pi}\left[\frac{\eta}{1-e^{-\eta}}\sum_{t=1}^{T}\widetilde{s}(\pi\prec\mu_{t})+\frac{\mathrm{KL}(\pi\|\pi_{0})}{1-e^{-\eta}}\right]$
@@ Appendix D More Related Work
Intransitivity in Game Theory. The symmetric zero-sum game and its intransitivity have also been frequently studied in the context of game theory. Balduzzi et al. (2018) was motivated by evaluation among different agents, showing that any symmetric zero-sum game can be decomposed into a “transitive” game and a “cyclic” game, and proposed Nash averaging for better agent/task evaluation. Balduzzi et al. (2019) generalized the results from matrix games to functional-form games and propose new algorithms to construct diverse populations of effective agents. Czarnecki et al. (2020) investigated the geometrical properties of real-world games (e.g., Tic-Tac-Toe, Go, StarCraft II) and proposed that real-world games have a “spinning top” geometry, with a strong transitive dimension and gradually diminishing non-transitive cyclic dimensions. Very recently, Bertrand et al. (2023) examined the limitations of the Elo rating system and proposed an alternative “disc decomposition” method that can better handle both transitive and cyclic game dynamics.
Representation Learning and Embedding. Representation learning and embedding techniques have successfully captured relational structures across various domains (Mikolov et al., 2013; Chen et al., 2020; Radford et al., 2021), yet their application in preference modeling and RLHF remains limited. Our work introduces preference representation learning, an approach that enhances expressiveness while maintaining computational efficiency, bridging the gap left by traditional approaches.
@@ Appendix E More on Experiments
Cyclic Preference Dataset. We constructed a dataset by inducing cyclic preferences from the Ultrafeedback dataset Cui et al. (2024). The dataset includes responses evaluated across four key metrics: instruction following, honesty, truthfulness, and helpfulness. We created preference cycles such as: instruction following $\succ$ honesty $\succ$ truthfulness $\succ$ helpfulness $\succ$ instruction following, ensuring the presence of intransitive cycles. We further generated four sub-datasets by omitting one metric from each cycle, resulting in 4 different datasets with 216 to 363 instances.
@@ E.2 Additional Experimental Results
Remark E.1.
We can have the following length-normalized (Meng et al., 2024) GPO (LN-GPO) learning objective (we report the empirical result in Table 7):
$\mathcal{L}_{\text{LN-GPO}}(\bm{\theta})=$ $\mathbb{E}_{\mathbf{x}\sim\mathcal{X},\mathbf{y}\sim\pi_{\bm{\theta}_{t}}(\cdot\mid\mathbf{x})}\Bigg{[}\Bigg{(}{\frac{1}{|y|}}\log\left(\frac{\pi_{\bm{\theta}}(\mathbf{y}\mid\mathbf{x})}{\pi_{\bm{\theta}_{t}}(\mathbf{y}\mid\mathbf{x})}\right)-\frac{1}{\beta}\left(\widehat{s}\left(\mathbf{y}\succ\pi_{\bm{\theta}_{t}}\mid\mathbf{x}\right)-\log Z_{\pi_{\bm{\theta}_{t}}}(\mathbf{x})\right)\Bigg{)}^{2}\Bigg{]}.$ (E.1)

### v3

@@ 2 Related Work
Reward-Based Reinforcement Learning from Human Feedback (RLHF). Typical approaches to modeling human preference for language model alignment usually learn a reward model from a preference dataset. The human preference is assumed to follow the Bradley-Terry (BT) model (Bradley & Terry, 1952) or the Thurstone model (Thurstone, 2017). LLM policies then are fine-tuned to maximize these scalar reward signals for better alignment (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022). Later, the direct preference optimization (DPO) methods are proposed by Rafailov et al. (2024) only implicitly to learn a reward model represented by an LLM. The human preference is still assumed to follow the Bradley-Terry model. However, the reliance on scalar rewards imposes a total ordering on preferences, which may not reflect the intransitive or stochastic nature of human judgments (Tversky, 1969; Agranov & Ortoleva, 2017).
Figure 2: Intransitiveness in real-world preferences. Left: Food preferences might cycle (Apple $\succ$ Banana $\succ$ Cherry $\succ$ Apple). Right: Rock-Paper-Scissors is a classic intransitive game.
Preference-Based Reinforcement Learning from Human Feedback. Recently, there emerged a line of works that directly estimate the preference probability without imposing a reward-based preference model or any transitivity assumptions (Lou et al., 2022; Wu et al., 2023; Wang et al., 2023) either for preference-based RL or in the context of RLHF. Efforts have been made to optimize policies directly from pair-wise preference comparisons, thereby mitigating the limitations of scalar reward functions (Munos et al., 2023; Swamy et al., 2024; Rosset et al., 2024; Wu et al., 2024b).
@@ 3.2 Preference Modeling
3.3 Pair Preference Models
@@ 3.4 Preference-based Reinforcement Learning from Human Feedback
$\max_{\pi}\ \mathbb{E}_{\mathbf{x}\sim\mathcal{X}}\bigg{[}$ $\mathbb{E}_{\mathbf{y}\sim\pi(\cdot\mid\mathbf{x}),\ \mathbf{y}^{\prime}\sim\mu(\cdot\mid\mathbf{x})}\Big{[}\mathbb{P}\big{(}\mathbf{y}\succ\mathbf{y}^{\prime}\mid\mathbf{x}\big{)}\Big{]}$
$-\beta\,\mathrm{KL}\big{(}\pi\|\pi_{\text{ref}}\big{)}\bigg{]},$
@@ 4.2 Implementing General Preference Embedding Model
To effectively capture general preferences while maintaining computational efficiency, we implement our preference embedding model by augmenting an existing language model with two additional components: an eigenvalue scale gate and an eigenvector embedding head. The embeddings $\mathbf{v}_{\mathbf{y}|\mathbf{x}}$ are derived from the final hidden state of the language model after processing the prompt $\mathbf{x}$ and response $\mathbf{y}$.
@@
where $\mathbf{D}(\mathbf{x})$ is a block-diagonal matrix with blocks $\sqrt{\lambda_{l}(\mathbf{x})}\mathbf{I}_{2}$, and $\mathbf{R}^{\succ}$ is the skew-symmetric preference operator. We normalize the embeddings $\mathbf{v}_{\mathbf{y}|\mathbf{x}}$ to have unit length $\|\mathbf{v}_{\mathbf{y}|\mathbf{x}}\|_{2}=1$ to ensure training stability and boundedness of the preference score (see Appendix C for details on boundedness related to Theorem 5.1).
@@ 5 General Preference Optimization
$\bm{\theta}_{t+1}=$ $\arg\min_{\bm{\theta}}\mathbb{E}_{\mathbf{x}\sim\mathcal{X},\mathbf{y}\sim\pi_{\bm{\theta}_{t}}(\cdot\mid\mathbf{x})}\Bigg{[}\bigg{(}\log\frac{\pi_{\bm{\theta}}(\mathbf{y}\mid\mathbf{x})}{\pi_{\bm{\theta}_{t}}(\mathbf{y}\mid\mathbf{x})}$
$-\frac{1}{\beta}\big{(}\widehat{s}\left(\mathbf{y}\succ\pi_{\bm{\theta}_{t}}\mid\mathbf{x}\right)-\log Z_{\pi_{\bm{\theta}_{t}}}(\mathbf{x})\big{)}\bigg{)}^{2}\Bigg{]},$ (5.3)
@@
Consider the optimization problem defined by the GPO loss (5) and assume it is realizable. Let $\{\pi_{\bm{\theta}_{t}}\}_{t=1}^{T}$ denote the sequence of policies generated by GPO, and define $\bar{\pi}_{T}=\frac{1}{T}\sum_{t=1}^{T}\pi_{\bm{\theta}_{t}}$ as the average policy. Given that the preference score $s$ is bounded within $[-\rho,\rho]$, by setting $\beta=\Theta\left(\sqrt{T}\right)$, we have:
@@
$\mathbb{E}_{\mathbf{x}\sim\mathcal{X},\mathbf{y}\sim\pi_{\bm{\theta}}}\bigg{[}\bigg{(}\widehat{s}(\mathbf{y}\succ\pi_{\bm{\theta}_{t}})-\beta\log\frac{\pi_{\bm{\theta}}(\mathbf{y}|\mathbf{x})}{\pi_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})}\bigg{)}$
$\hskip 86.11084pt\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(\mathbf{y}|\mathbf{x})\bigg{]}$
@@
So Equation (5) can also be seen as a policy gradient method for the optimization problem (5.1).
@@ 6 Experiments
Table 3: AlpacaEval 2.0 evaluation results. Base model: Llama3-8B-it. Evaluator: GPT-4-turbo. Results grouped by RM/GPM size (2B/8B) and type (BT/GPM), and training iterations (Iter). Bold entries indicate GPM outperforms BT RM under the same SPPO/GPO setting. Metrics follow AlpacaEval 2.0 standard: LC. WR = Length-Controlled Win Rate (%), WR = Win Rate (%) against the reference model (GPT-4-turbo), Avg. Len = Average response length (tokens).
@@ 6.1 Experiments on RewardBench
Results and Analysis. The results are presented in Table 2. On RewardBench, using the Gemma-2B-it base model, GPM achieves an average score of 82.29%, which is an improvement of 7.44% over the BT model’s average score of 74.85%. Specifically, in the Chat task, GPM improves performance from 67.32% (BT RM) to 79.61%, and in the Chat-Hard task, from 63.37% to 75.66%. For the Llama-3.1-8B-Instruct base model, GPM achieves an average score of 91.90% (embedding dimension 8), representing a 1.34% improvement over the BT model’s average score of 90.56%. In the Chat task, GPM improves from 88.55% (BT RM) to 93.58%, and in the Chat-Hard task, from 85.75% to 88.38%. These results indicate that GPM generally outperforms the BT model across various base models and tasks, particularly in the Chat and Chat-Hard categories which may involve more nuanced or complex preferences. Note that BT RM is a special case of GPM when the embedding dimension $d=1$ (see Section 4.3). Further analysis on challenging cases is in Appendix E.1.
@@ 6.3 Downstream Performance on Aligning Language Models with Human Preferences
We further investigate the effectiveness of GPM in language model alignment using Self-Play Policy Optimization (SPPO) (Wu et al., 2024b) and our proposed General Preference Optimization (GPO), integrating preference scores provided by our GP representation model (GPM). We evaluated the models on AlpacaEval 2.0 (Dubois et al., 2024), MT-Bench (Zheng et al., 2023), GSM8K, MMLU, etc., several widely used benchmarks for evaluating LLM alignment.
Results and Analysis. The evaluation results on the benchmarks are as follows. For AlpacaEval 2.0, we compared the generated responses of the aligned models with those of GPT-4o-mini and GPT-4-turbo. The results of the three evaluators are presented in Tables 3 and 5. From Table 3, we observe that both SPPO and GPO demonstrate improved win rates with successive iterations, highlighting the iterative nature of these optimization methods, and GPO consistently outperforms SPPO. In addition, the bolded entries indicate that GPM-integrated methods consistently outperform BT RM-based methods under the same settings on Win Rate (WR).
@@ 7 Conclusion
Impact Statement
This paper presents work to align AI systems with human preferences and values. Improving AI alignment could lead to more reliable and helpful AI systems that better serve human needs while respecting human values. The enhanced ability to model complex, intransitive preferences could result in AI systems that better understand and accommodate nuanced human judgments across different contexts. This advancement underscores the potential of AI alignment algorithms in both technological and societal contexts.
Acknowledgements
We thank Andrew Yao, Yang Yuan, Kaifeng Lyu, Haodong Wen, and Yifan Luo for helpful discussions and feedback. We also thank the anonymous reviewers for their valuable comments and suggestions, which helped us significantly to improve this paper.
@@ 2. Magnitude preserving:
The proof follows the logic of standard multiplicative weights update analysis, adapted for our preference score objective. First, since the preference score $s$ is bounded in $[-\rho,\rho]$, we can normalize it to $[0,1]$ by the transformation:
@@
By Theorem 1 in (Freund & Schapire, 1999), for any sequence of mixed policies $\mu_{1},\mu_{2},\ldots,\mu_{T}$, the sequence of policies $\pi_{1},\pi_{2},\ldots,\pi_{T}$ produced by GPO satisfies:
$\sum_{t=1}^{T}\widetilde{s}(\pi_{t}\prec\mu_{t})\leq\min_{\pi}\left[\frac{\eta}{1-e^{-\eta}}\sum_{t=1}^{T}\widetilde{s}(\pi\prec\mu_{t})+\frac{\mathrm{KL}(\pi\|\pi_{0})}{1-e^{-\eta}}\right],$
where $\eta=1/\beta$ is the learning rate parameter in the MWU view (related to $\beta$ in our GPO loss Eq. 5). The preference score $s$ is assumed bounded, which holds in our implementation due to L2 normalization of embeddings and bounded eigenvalues (Section 4.2).
@@ Appendix D More Related Work
Intransitivity in Game Theory. The symmetric zero-sum game and its intransitivity have also been frequently studied in the context of game theory. Balduzzi et al. (2018) explored game decompositions into transitive and cyclic components, proposing Nash averaging for agent evaluation. Balduzzi et al. (2019) generalized the results from matrix games to  functional-form games and proposed algorithms for diverse agent populations. Czarnecki et al. (2020) investigated the geometrical properties of real-world games (e.g., Tic-Tac-Toe, Go, StarCraft II) and proposed that real-world games have a “spinning top” geometry, with a strong transitive dimension and gradually diminishing non-transitive cyclic dimensions. Very recently, Bertrand et al. (2023) examined the limitations of the Elo rating system and proposed an alternative “disc decomposition” method that can better handle both transitive and cyclic game dynamics.
Representation Learning and Embedding. Representation learning involves learning transformations of data that make it easier to extract useful information (Bengio et al., 2013). Techniques like word embeddings (Mikolov et al., 2013) and contrastive learning (Chen et al., 2020; Radford et al., 2021) learn powerful representations for various tasks. While these methods capture semantics or similarities, their direct application to modeling the directed, potentially asymmetric, and intransitive nature of human preferences in RLHF has been less explored. Our GPM specifically introduces a structured embedding space with a skew-symmetric operator tailored to capture these preference characteristics, offering a novel application of representation learning principles to the alignment problem.
Alternative Preference/Reward Models. Beyond the standard BT model, other approaches exist. Pairwise preference models (Section 3.3) often concatenate pairs, potentially suffering from order effects and quadratic complexity in the number of responses per prompt. Multi-dimensional reward models like ArmoRM (Wang et al., 2024a) decompose reward into interpretable dimensions (e.g., helpfulness, harmlessness). While offering interpretability, they typically require multi-dimensional annotations and, being based on summing scalar rewards, may still struggle to represent complex intransitive structures compared to GPM’s more general formulation (Theorem 4.4). GPM aims for expressiveness and automatic discovery of preference dimensions from standard pairwise labels.
Preference Optimization Methods. Recent work has explored various algorithms beyond PPO for RLHF. DPO (Rafailov et al., 2024) directly optimizes the policy using a derived reward. IPO (Azar et al., 2023) and related works (Mitchell, 2023; Liang et al., 2024; Furuta et al., 2024) focus on robustness or handling noisy/soft preferences within the DPO framework. For instance, Mitchell (2023) analyzed the relationship between DPO and IPO under noisy preferences, while Liang et al. (2024) and Furuta et al. (2024) proposed robust preference optimization methods to handle noise and soft labels, respectively. Nash-based methods (Munos et al., 2023; Swamy et al., 2024; Rosset et al., 2024; Wu et al., 2024b) explicitly consider the game-theoretic aspect of finding optimal policies. Our GPO (Section 5) builds upon these ideas but uses the preference score $s$ directly, offering an alternative optimization target compared to win rates or implicit rewards. GPM’s preference scores can potentially be integrated into many of these existing optimization frameworks.
@@ Appendix E More on Experiments
Cyclic Preference Dataset. We constructed a dataset by inducing cyclic preferences from the Ultrafeedback dataset (Cui et al., 2024). The dataset includes responses evaluated across four key metrics: instruction following, honesty, truthfulness, and helpfulness. We created preference cycles such as: instruction following $\succ$ honesty $\succ$ truthfulness $\succ$ helpfulness $\succ$ instruction following, ensuring the presence of intransitive cycles. We further generated four sub-datasets by omitting one metric from each cycle, resulting in 4 different datasets with 216 to 363 instances.
We construct the Cyclic Preference Dataset using the following steps: 1. Filter Ultrafeedback for prompts with at least 3 responses where multi-aspect ratings (helpfulness, honesty, etc.) are available.
2. For a set of 3 responses (A, B, C) to the same prompt, check if ratings imply a cycle based on chosen criteria. Example: If criterion 1 is Honesty and criterion 2 is Helpfulness, check if Rating(A, Honesty) $\succ$ Rating(B, Honesty), Rating(B, Helpfulness) $\succ$ Rating(C, Helpfulness), and Rating(C, Honesty) $\succ$ Rating(A, Honesty). This implies a potential cycle A $\succ$ B $\succ$ C $\succ$ A across criteria. 3. Define pairwise preferences based on the higher rating for the specific criterion involved in the cycle link (e.g., A $\succ$ B based on Honesty, B $\succ$ C based on Helpfulness, C $\succ$ A based on Honesty). 4. Collect these cyclic triplets (or longer cycles) to form the datasets used in Table 1. Different datasets (Cyclic No. 1-4) were created by focusing on cycles involving different combinations of the four criteria.
@@ E.2 Additional Experimental Results
Discussion on Length Control (LC. WR). As observed in Section 6.3 (Table 3, 5), GPM-aligned models often produce longer responses than BT-aligned models. The standard AlpacaEval Win Rate (WR) doesn’t penalize length, whereas the Length-Controlled Win Rate (LC. WR) is designed to mitigate length bias, potentially penalizing models that win primarily by being more verbose. GPM’s multi-dimensional embeddings might better capture the value of comprehensive answers that address multiple facets of a prompt (e.g., being helpful, detailed, and stylistically appropriate simultaneously), leading to higher preference scores and thus longer generated responses under GPO. While beneficial for overall quality perceived by the preference model, this can negatively impact the LC. WR metric. For applications where conciseness is highly valued or length bias is a major concern, specific techniques might be needed. One approach is length normalization within the optimization objective.
Length-Normalized GPO (LN-GPO) Results. We explored a variant, Length-Normalized GPO (LN-GPO), which modifies the GPO loss (Eq. 5) to normalize the policy update by response length, similar in spirit to Meng et al. (2024). The objective becomes:
$\mathcal{L}_{\text{LN-GPO}}(\bm{\theta})=$ $\mathbb{E}_{\mathbf{x}\sim\mathcal{X},\mathbf{y}\sim\pi_{\bm{\theta}_{t}}(\cdot\mid\mathbf{x})}\Bigg{[}\Bigg{(}{\frac{1}{|\mathbf{y}|^{\gamma}}}\log\left(\frac{\pi_{\bm{\theta}}(\mathbf{y}\mid\mathbf{x})}{\pi_{\bm{\theta}_{t}}(\mathbf{y}\mid\mathbf{x})}\right)-\frac{1}{\beta}\left(\widehat{s}\left(\mathbf{y}\succ\pi_{\bm{\theta}_{t}}\mid\mathbf{x}\right)-\log Z_{\pi_{\bm{\theta}_{t}}}(\mathbf{x})\right)\Bigg{)}^{2}\Bigg{]}.$
Here, $|\mathbf{y}|$ is the length of the response $\mathbf{y}$, and $\gamma$ is a hyperparameter (typically 1). Initial results using LN-GPO with $\gamma=1$ are shown in Table 7. We observe that LN-GPO with GPM (2B) achieves a slightly higher LC. WR (45.55%) compared to LN-GPO with BT RM (45.51%), while producing longer responses (Avg. Length 2112 vs 1951). This suggests length normalization can help, although further tuning and investigation are needed to fully balance performance and length control with GPM.